OpenStudy (anonymous):

Use implicit differentiation to find dy/dx where x^2y=1 + e^y

OpenStudy (anonymous):

Do you know how to do Implicit D?

OpenStudy (anonymous):

A little bit. But I'm confused with the e^y

OpenStudy (anonymous):

Chain rule it. \[\left(e^{u(x)}\right )^{\prime} = e^{u(x)} (u(x))^{\prime}\] right? just replace u(x) with y(x)

OpenStudy (anonymous):

I have never seen that before... I'm familiar with the chain rule but have never seen it done like that.

OpenStudy (anonymous):

So they probably taught you the chain rule by \[\frac{d}{dx}\bigg(stuff(x)\bigg)^7= 7\bigg(stuff(x)\bigg)^6\] or something like that?

OpenStudy (anonymous):

yeah exactly

OpenStudy (anonymous):

Yeah its easier to understand what to do that way but harder on explaining Implicit D. Notice that I actually finished wrong it should have been \[7\bigg(stuff(x)\bigg)^6 \bigg(stuff(x) \bigg)^{\prime}\] right? the issue here is that you are taught basic derivatives like \[ f(x) = x^4\to f^{\prime}(x) = 4x^3\] So what happens when you have \[f(x) = (g(x))^4\] It looks like \[ f(x) = x^4\to f^{\prime}(x) = 4x^3\] but who knows what g(x) is. So they came up with this cool fact that \[\frac{df}{dx}=\frac{df}{dg}\frac{dg}{dx}\] Notice that \[\frac{df}{dg} = 4(g(x))^3\] and \[\frac{dg}{dx}= (g(x))^{\prime}\] It basically mimics the stuff argument but with a variable for "stuff(x)" So instead of actually finding the (g(x))' you leave it for implicit D and then factor and solve for it.

OpenStudy (anonymous):

I understand it a little bit more, but I'm still so confused.

OpenStudy (anonymous):

Do you know of any good websites or books that can help me with that? My textbook is very confusing.

OpenStudy (anonymous):

just math tutoring has a video on this that makes it look painfully easy the explanation is something that will make a math teacher cringe, but it works i will see if i can find it

OpenStudy (anonymous):

unfortunately i cannot link to the video exactly but it is here, just scroll down to "calculus" and then "implicit differentiation" http://justmathtutoring.com/

OpenStudy (anonymous):

here's a PDF I wrote. It gives an example and reviews the chain rule.

OpenStudy (anonymous):

justmathtutoring aka patrickjmt.com? yeah i use his site a lot, but i didn't see a problem with e^y

OpenStudy (anonymous):

okay thanks guys. i'll keep at it.

OpenStudy (anonymous):

Remember that e^y is just a function that takes y as a variable. you could argue that its e(y) or as it is sometimes written exp(y).

OpenStudy (anonymous):

just like cos(y)